OpenStudy (ibeyoshiii):

If g(x) = 3(x − 2), find the value of x if g(x) = 6.

OpenStudy (anonymous):

12

OpenStudy (anonymous):

basically, plug in 6 where you see an x

OpenStudy (ibeyoshiii):

Oh. Okay. Cool. Thanks.

OpenStudy (anonymous):

g(x) = 3(x − 2) g(6) = 3(6-2) g(6) = 3(4) g(6) = 12

OpenStudy (ibeyoshiii):

Thanks. :D